Molar mass of Glucose is:

(A) 180 g
(B) 129 g
(C) 190 g
(D) 200 g

Grade:12th pass

3 Answers

Vikas TU
3 years ago
Dear studnet 
a mole is the quantity of a substance whose weight in grams is equal to the molecular weight of the substance. Thus 1 mole of glucose weighs 180 g.
Vikas Amritiya
3 years ago

 

Glucose is a simple sugar with the molecular formula C6H12O6.It is soluble in water and acetic acid.
 
Molar mass of glucose can be calculated by the sum of relative atomic masses of atoms in a molecule of a substance.
 
As we know, atomic mass of C = 12
                      atomic mass of H = 1
                      atomic mass of O = 16
 
Molar mass of C6H12O6 = 6×12+1×12+6×16 = 72+12+96 = 180
 
So, the molar mass of glucose is 180g
 
(A)  180g is correct answer.
irshad hussain
2 years ago

Let us look for the atomic masses of carbon (C), hydrogen (H) and oxygen (O).

C = 12

H = 1

O = 16

∴ The molar mass of the glucose is

⇒ 12 x 6 + 1 x 12 + 16 x 6

⇒ 72 + 12 + 96

= 180